# HMSViT: A Hierarchical Masked Self-Supervised Vision Transformer for Corneal Nerve Segmentation and Diabetic Neuropathy Diagnosis

### v2

@@ Abstract
Diabetic Peripheral Neuropathy (DPN) affects nearly half of diabetes patients, requiring early detection. Corneal Confocal Microscopy (CCM) enables non-invasive diagnosis, but automated methods suffer from inefficient feature extraction, reliance on handcrafted priors, and data limitations. We propose HMSViT, a novel Hierarchical Masked Self-Supervised Vision Transformer (HMSViT) designed for corneal nerve segmentation and DPN diagnosis. Unlike existing methods, HMSViT employs pooling-based hierarchical and dual attention mechanisms with absolute positional encoding, enabling efficient multi-scale feature extraction by capturing fine-grained local details in early layers and integrating global context in deeper layers, all at a lower computational cost. A block-masked self supervised learning framework is designed for the HMSViT that reduces reliance on labelled data, enhancing feature robustness, while a multi-scale decoder is used for segmentation and classification by fusing hierarchical features. Experiments on clinical CCM datasets showed HMSViT achieves state-of-the-art performance, with 61.34% mIoU for nerve segmentation and 70.40% diagnostic accuracy, outperforming leading hierarchical models like the Swin Transformer and HiViT by margins of up to 6.39% in segmentation accuracy—while using fewer parameters. Detailed ablation studies further reveal that integrating block-masked SSL with hierarchical multi-scale feature extraction substantially enhances performance compared to conventional supervised training. Overall, these comprehensive experiments confirm that HMSViT delivers excellent, robust, and clinically viable results, demonstrating its potential for scalable deployment in real-world diagnostic applications.
@@
Diabetic Peripheral Neuropathy (DPN), Corneal Confocal Microscopy (CCM), Vision Transformer, Self-Supervised Learning (SSL), Hierarchical Multi-Scale Learning
@@ 1 Introduction
[…]
Diabetic Peripheral Neuropathy (DPN) is a debilitating complication of diabetes, affecting nearly 50% of individuals with diabetes worldwide and significantly increasing the risk of foot ulcers, infections, and lower-limb amputations [1]. This progressive neurodegenerative condition is characterized by nerve damage, primarily in the peripheral nervous system, leading to symptoms such as pain, numbness, and loss of sensation in the extremities. Early diagnosis and continuous monitoring are essential to mitigate its impact and prevent irreversible nerve damage.
@@
Nonetheless, standard ViTs face notable challenges in medical imaging, especially for tasks demanding high-resolution representations, such as organ segmentation or nerve fibre analysis [9, 10]. Firtstly, ViTs require extensive annotated datasets, which are often unavailable in medical contexts, and their patch-based processing reduces spatial resolution and lacks inductive bias for spatial reasoning - the ability to understand structure and relationships within an image. These limitations call for innovative solutions to harness ViTs’ strengths while addressing their shortcomings.
@@
To address these issues, we propose HMSViT—a Hierarchical Masked Self-Supervised Vision Transformer. HMSViT employs a pooling-based hierarchical design with dual attention mechanisms, where early layers capture fine-grained local details and deeper layers integrate global contextual information using absolute positional encodings. A novel block-masked SSL approach is introduced to learn robust spatial relationships from unlabelled CCM images, reducing reliance on extensive annotations. A multi-scale decoder further fuses these features for both accurate corneal nerve segmentation and DPN diagnosis. In summary, our contributions are threefold:
@@
    Efficient Multi-Scale Feature Extraction: A pooling-based hierarchical design with dual attention effectively captures both local and global features without heavy reliance on handcrafted modules.
@@
    Innovative Self-Supervised Learning: A block-masked SSL approach tailored for hierarchical architectures improves spatial reasoning while minimizing computational overhead.
@@
    State-of-the-Art Clinical Performance: HMSViT demonstrates superior segmentation accuracy and diagnostic classification on clinical CCM datasets, offering a scalable, non-invasive solution for early DPN screening.
@@ 2 Related work
2.1 Epidemiological Context and Clinical Need
Diabetes affects 463 million people worldwide, projected to reach 700 million by 2045 [13]. Diabetic Peripheral Neuropathy (DPN), impacting up to 50% of patients, causes 50–75% of non-traumatic lower-limb amputations [14]. Defined as peripheral nerve dysfunction in diabetes after excluding other causes [15], DPN highlights the urgent need for early detection to prevent irreversible damage.
Corneal Confocal Microscopy (CCM) is a non-invasive tool for early DPN detection, allowing quantitative analysis of the corneal sub-basal nerve plexus, which is highly sensitive to metabolic damage [16]. Manual annotation of CCM images is time-consuming, expertise-dependent, and subject to variability due to noise and interpretation bias [5]. These limitations have led to the development of automated methods to extract key diagnostic metrics, including corneal nerve fibre length (CNFL), density (CNFD), and branch density (CNBD).
@@ 2.2 Deep Learning Methods for Corneal Nerve Segmentation and DPN Diagnosis
Recent advancements in deep learning have significantly improved automated CCM image analysis, particularly in the areas of corneal nerve segmentation and DPN classification. The most prominent approaches include Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs), both of which have demonstrated success in various medical imaging applications.
@@ 2.2.1 CNN-Based Approaches
CNNs have been extensively used for DPN classification and corneal nerve segmentation, leveraging their ability to capture hierarchical spatial features. Several CNN-based approaches have been explored in recent years. [17] developed a CNN-based classifier for DPN detection using CCM images, incorporating data augmentation techniques to improve model generalisation. Similarly, [18] applied ResNet architectures, achieving high classification accuracy in identifying neuropathic changes. For segmentation, [19] employed U2Net to precisely quantify corneal nerve fibre parameters, significantly enhancing the sensitivity of DPN detection. [7] further demonstrated that CNN-based segmentation approaches could outperform traditional corneal nerve analysis software (e.g., ACC Metrics), providing more accurate and reliable quantification of nerve fibre morphology. In addition, [20] implemented a 3D CNN model to classify painful DPN subtypes, highlighting the potential of CNNs in personalized medicine and treatment response analysis.
Despite their effectiveness, CNN-based approaches are fundamentally constrained by fixed local receptive fields, limiting their ability to model long-range dependencies and complex nerve fibre branching patterns. Since CNNs rely on localized feature extraction, they struggle to capture the global structure of corneal nerve morphology, which is critical for accurate segmentation. Additionally, CNN-based models are highly sensitive to image contrast variations and noise, making their performance inconsistent across different datasets. These limitations have led researchers to explore Vision Transformers (ViTs), which leverage self-attention mechanisms to capture both local and global spatial dependencies.
@@ 2.2.2 Vision Transformers (ViTs)
Standard Vision Transformer (ViT) Hierarchical Vision Transformer (HVT) Patch Handling Fixed-size patches across all layers Progressive patch merging & down sampling Computate Complexity Quadratic in image resolution Reduced complexity with hierarchical tokenisation Feature Extraction Uniform token processing Multi-scale feature learning (local to global) Examples Original ViT[8], DeiT [23], DPT [24] Swin Transformer[11], PVT [25], HiVi[12] Performance Performs well on image classification, especially with large datasets and pretraining. Often achieves better performance on dense prediction tasks like object detection and semantic segmentation due to multi-scale feature representation.
Table 1: Comparison of Standard Vision Transformer (ViT) and Hierarchical Vision Transformer (HVT) architectures.
To address the limitations of standard ViTs, Hierarchical Vision Transformers (HVTs) introduce multi-scale feature extraction, progressively refining spatial resolution through patch merging and downsampling. Table 1 provides a concise comparison between standard ViTs and Hierarchical ViTs. Standard ViTs process images with fixed-size patches across all layers, resulting in uniform tokenisation . While effective, this approach leads to quadratic computational complexity and requires large datasets for training. In contrast, Hierarchical ViTs use progressive patch merging and downsampling. This reduces computational cost and enables multi-scale feature learning. As a result, HVTs capture both local and global structures efficiently and often generalise better on smaller datasets by leveraging spatial hierarchies.
@@
Despite these promising developments, the application of hierarchical vision transformers (HVTs) in medical imaging remains challenging. Although HVTs offer multi-scale representation, they typically depend on spatial priors such as convolutions [12], shifted windows [11], and intricate relative position embeddings to introduce spatial bias into the model [27]. While these mechanisms improve supervised task accuracy, they also increase computational complexity, reduce inference speed, and complicate the integration of efficient self-supervised tasks.
Moreover, no existing studies have explored the application of HVTs for corneal confocal microscopy (CCM) image segmentation or DPN diagnosis, leaving a critical gap in leveraging their potential for medical imaging tasks. Additionally, the reliance on large-scale annotated datasets for effective pretraining presents a major obstacle in medical applications, where labelled data is often scarce, costly, and time-consuming to obtain. To address these limitations, recent efforts have focused on integrating self-supervised learning (SSL) techniques with vision transformer architectures. SSL enables models to learn robust spatial representations from abundant unannotated data, significantly reducing their dependency on manual annotations and improving their adaptability to medical imaging tasks.
@@ 2.3 Self-Supervised Learning (SSL)
Self-supervised learning (SSL) has become essential in computer vision, enabling models to learn robust representations from large unlabelled datasets. This approach is particularly valuable for Vision Transformers (ViTs), which require extensive data to achieve competitive performance compared to convolutional neural networks [28]. SSL methods rely on pretext tasks such as predicting missing image parts or distinguishing between different augmented views of an image [29], helping models capture spatial and semantic information without labelled data.
@@
In contrast, we present block masked self-supervised learning for hierarchical ViT pretraining, resulting in a model that is both powerful and efficient.
@@ 3 The Proposed Method
    Hierarchical Multi-Scale Feature Extraction: HMSViT processes images through four stages, starting with high-resolution inputs to capture fine details and gradually reducing spatial resolution to extract higher-level features. Instead of relying on complex modules (e.g., shifted windows or convolutions), our design uses a non-parametric pooling operator, with block-based local attention in early stages and global attention in deeper stages complemented by block-level absolute positional encodings.
@@ 3.1 Hierarchical Multi-Scale Vision Transformer (HMSViT)
The proposed method leverages the strengths of Vision Transformers (ViTs) while addressing their limitations in dense medical imaging tasks—specifically, for corneal nerve fibre segmentation and Diabetic Peripheral Neuropathy (DPN) diagnosis. By progressively extracting features at multiple scales, the network captures both fine-grained local details and broader global context, leading to enhanced segmentation performance and improved diagnostic accuracy.
@@ 3.1.1 High-Resolution Hierarchical Representations
where $L=\frac{H}{p}\times\frac{W}{p}$ is the number of patches, $p$ is the patch size.$d$ is the feature dimension. $\frac{H}{p}\times\frac{W}{p}$ is the output feature size.
@@
A non-parametric max pooling operator is then applied to each block to downsample the representation while preserving the most salient features. The pooling operation is defined as:
@@ 3.1.2 Dual Attention Mechanisms
Given that high-resolution features in the early stages result in $L=\frac{H}{p}\times\frac{W}{p},p=4$, applying global self-attention becomes computationally intensive due to its O(L²) complexity. To mitigate this, HMSViT employs a dual attention strategy tailored to different stages of the network.
$\text{Attention}=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)V$ (4)
  • Block-based Attention in early stages, where attention is computed within local blocks of 16 patches.
  • Global Attention in deeper stages, as the spatial dimensions reduce, the model switches to global attention to integrate high-level semantic information across the entire image.
@@ 3.1.3 Absolute Positional encodings
To preserve spatial relationships—crucial in medical imaging—we add positional encodings to the input data. While standard ViTs use absolute or relative embeddings (with hierarchical models often favoring relative ones for consistency across varying resolutions), relative embeddings require modifying the attention mechanism and add computational overhead. In our method, we simply add learnable absolute positional encodings at the block level in deeper stages:
@@
3.2 Block Masked self-supervised learning
Figure 2: Block Masked Self Supervised Learning
Our approach builds on the recent advances in Masked Autoencoders (MAE) [36], where a large portion of the input (typically 75–90%) is masked out, and the model is trained to reconstruct the missing parts. This strategy forces the network to learn rich spatial relationships, as it must infer the structure of the masked regions from the visible context. However, in a hierarchical Vision Transformer (ViT) with progressive downsampling, directly applying standard MAE can disrupt spatial coherence across scales. To address this, we propose a block-masking strategy that groups small patches (e.g., 4×4 patches) into larger block units before masking. As shown in Figure 2, we (1) group the patches into block units, (2) randomly mask blocks at the block level, and (3) ungroup them back into smaller patches for our HMSViT.

### v3

# HMSViT: A Hierarchical Masked Self-Supervised Vision Transformer for Efficient and Robust Corneal Nerve Segmentation and Diabetic Neuropathy Diagnosis

@@ Abstract
Diabetic peripheral neuropathy (DPN) is a common and disabling complication of diabetes, necessitating early and accurate detection. Corneal confocal microscopy (CCM) offers a non-invasive diagnostic window, but automated analysis is often limited by inefficient feature extraction and scarce annotations. We propose HMSViT, a hierarchical masked self-supervised vision transformer for corneal nerve segmentation and DPN diagnosis. HMSViT uses a pooling-based hierarchical architecture with a dual-attention design and absolute positional encoding to capture multi-scale features efficiently. The model is pretrained with block-masked self-supervised learning to improve robustness and reduce reliance on labelled data, and a lightweight multi-scale decoder is used for downstream segmentation and classification. Experiments on clinical CCM datasets using patient-level 5-fold cross-validation demonstrate strong performance, achieving 85.6% patient-level diagnostic accuracy and 61.34% mIoU for nerve segmentation. Compared with Swin Transformer and HiViT, HMSViT improves segmentation mIoU by 2.45–3.04% and diagnostic accuracy by 2.3–3.8% while using up to 41% fewer parameters. Ablation studies show that combining hierarchical multi-scale design with block-masked pretraining is critical for optimal performance.
@@
Diabetic peripheral neuropathy (DPN), corneal confocal microscopy (CCM), self-supervised learning (SSL), hierarchical multi-scale learning.
@@ 1 Introduction
Diabetic peripheral neuropathy (DPN) is a debilitating complication of diabetes, affecting nearly 50% of individuals with diabetes worldwide and significantly increasing the risk of foot ulcers, infections, and lower-limb amputations [1]. This progressive neurodegenerative condition is characterised by nerve damage, primarily in the peripheral nervous system, leading to symptoms such as pain, numbness, and loss of sensation in the extremities. Early diagnosis and continuous monitoring are essential to mitigate its impact and prevent irreversible nerve damage.
@@
Nonetheless, standard ViTs face notable challenges in medical imaging, especially for tasks demanding high-resolution representations, such as organ segmentation or nerve fibre analysis [9, 10]. Firstly, ViTs require extensive annotated datasets, which are often unavailable in medical contexts, and their patch-based processing reduces spatial resolution and lacks inductive bias for spatial reasoning - the ability to understand structure and relationships within an image. These limitations call for innovative solutions to harness ViTs’ strengths while addressing their shortcomings.
@@
To address these issues, we propose HMSViT—a Hierarchical Masked Self-Supervised Vision Transformer. Unlike existing hierarchical ViTs that rely on shifted windows or convolutional priors, HMSViT introduces a pooling-based hierarchical design with dual attention mechanisms, where early layers capture fine-grained local details and deeper layers integrate global contextual information using absolute positional encodings. A novel block-masked SSL approach is introduced to learn robust spatial relationships from unlabelled CCM images, reducing reliance on extensive annotations. A multi-scale decoder further fuses these features for both accurate corneal nerve segmentation and DPN diagnosis. In summary, the key contributions are:
@@
    We propose HMSViT, a novel hierarchical Vision Transformer architecture designed specifically for medical image analysis, which integrates dual attention mechanisms and pooling-based token aggregation to efficiently capture both fine-grained local details and global contextual information from high-resolution CCM images.
@@
    We introduce a novel block-masked self-supervised learning strategy tailored for hierarchical vision transformers, enabling robust feature representation from unlabelled data and significantly reducing reliance on extensive expert annotations—a key barrier in clinical AI deployment.
@@
    We demonstrate that HMSViT achieves state-of-the-art performance on real-world clinical datasets, outperforming leading models such as Swin Transformer and HiViT in both segmentation and diagnosis, while using fewer parameters and offering strong potential for scalable clinical deployment.
@@ 2.1 Deep Learning Methods for Corneal Nerve Segmentation and DPN Diagnosis
Diabetes affects 463 million people worldwide, projected to reach 700 million by 2045 [13]. DPN, impacting up to 50% of patients, causes 50–75% of non-traumatic lower-limb amputations [14]. Defined as peripheral nerve dysfunction in diabetes after excluding other causes [15], DPN highlights the urgent need for early detection to prevent irreversible damage.
Corneal Confocal Microscopy (CCM) is a non-invasive tool for early DPN detection, allowing quantitative analysis of the corneal sub-basal nerve plexus, which is highly sensitive to metabolic damage [16]. Manual annotation of CCM images is time-consuming, expertise-dependent, and subject to variability due to noise and interpretation bias [5]. These limitations have led to the development of automated methods to extract key diagnostic metrics, including corneal nerve fibre length (CNFL), density (CNFD), and branch density (CNBD). Recent advancements in deep learning have significantly improved automated CCM image analysis, particularly in the areas of corneal nerve segmentation and DPN classification. The most prominent approaches include Convolutional Neural Networks (CNNs) and ViTs, both of which have demonstrated success in various medical imaging applications.
@@ 2.1.1 CNN-Based Approaches
CNNs have been extensively used for DPN classification and corneal nerve segmentation, leveraging their ability to capture hierarchical spatial features. Several CNN-based approaches have been explored in recent years. [17] developed a CNN-based classifier for DPN detection using CCM images, incorporating data augmentation techniques to improve model generalisation. Similarly, [18] applied ResNet architectures, achieving high classification accuracy in identifying neuropathic changes. For segmentation, [19] employed U2Net to precisely quantify corneal nerve fibre parameters, significantly enhancing the sensitivity of DPN detection. [7] further demonstrated that CNN-based segmentation approaches could outperform traditional corneal nerve analysis software (e.g., ACC Metrics), providing more accurate and reliable quantification of nerve fibre morphology. In addition, [20] implemented a 3D CNN model to classify painful DPN subtypes, highlighting the potential of CNNs in personalised medicine and treatment response analysis.
Despite their effectiveness, CNN-based approaches are fundamentally constrained by fixed local receptive fields, limiting their ability to model long-range dependencies and complex nerve fibre branching patterns. Since CNNs rely on localized feature extraction, they struggle to capture the global structure of corneal nerve morphology, which is critical for accurate segmentation. Additionally, CNN-based models are highly sensitive to image contrast variations and noise, making their performance inconsistent across different datasets. These limitations have led researchers to explore ViTs, which leverage self-attention mechanisms to capture both local and global spatial dependencies.
@@ 2.1.2 Vision Transformers (ViTs)
Standard Vision Transformer (ViT) Hierarchical Vision Transformer (HVT) Patch Handling Fixed-size patches across all layers Progressive patch merging & downsampling Computational Complexity Quadratic in image resolution Reduced complexity with hierarchical tokenisation Feature Extraction Uniform token processing Multi-scale feature learning (local to global) Examples Original ViT[8], DeiT [23], DPT [24] Swin Transformer[11], PVT [25], HiViT[12] Performance Performs well on image classification, especially with large datasets and pretraining. Often achieves better performance on dense prediction tasks like object detection and semantic segmentation due to multi-scale feature representation.
Table 1: Comparison between Standard Vision Transformer (ViT) and Hierarchical Vision Transformer (HVT) architectures, highlighting differences in patch handling, computational complexity, feature extraction, representative models, and typical performance characteristics.
To address the limitations of standard ViTs, Hierarchical Vision Transformers (HVTs) introduce multi-scale feature extraction, progressively refining spatial resolution through patch merging and downsampling. Table 1 provides a concise comparison between standard ViTs and Hierarchical ViTs. Standard ViTs process images with fixed-size patches across all layers, resulting in uniform tokenisation. While effective, this approach leads to quadratic computational complexity and requires large datasets for training. In contrast, Hierarchical ViTs use progressive patch merging and downsampling, reducing computational cost and enabling multi-scale feature learning. As a result, HVTs efficiently capture both local and global structures and often generalise better on smaller datasets by leveraging spatial hierarchies (see Table 1).
@@
Despite these advances, the application of hierarchical HVTs in medical imaging remains challenging. Although existing HVTs offer multi-scale representations, they often rely on explicit spatial priors—such as convolutions [12], shifted windows [11], and complex relative positional encodings to encode locality and spatial bias. While these mechanisms improve supervised accuracy, they also increase computational complexity, slow inference, and reduce architectural flexibility. These drawbacks are particularly problematic in medical contexts, where efficiency and adaptability are critical. Moreover, no prior study has investigated the application of HVTs for CCM image segmentation or DPN diagnosis, leaving a significant gap in clinical applications.
At the same time, the scarcity of annotated datasets limits the effectiveness of purely supervised approaches in medical imaging. This has motivated the adoption of self-supervised learning (SSL), which allows models to exploit abundant unlabelled data to learn transferable representations. Yet, existing hierarchical designs are not well suited to SSL, as their reliance on heavy spatial priors complicates pretraining and reduces adaptability. These challenges highlight the need for new hierarchical methods that are both computationally efficient and naturally compatible with self-supervised learning.
@@ 2.2 Self-Supervised Learning (SSL)
Self-supervised learning (SSL) has become essential in computer vision, enabling models to learn robust representations from large unlabelled datasets. This approach is particularly valuable for ViTs, which require extensive data to achieve competitive performance compared to convolutional neural networks [27]. SSL methods rely on pretext tasks such as predicting missing image parts or distinguishing between different augmented views of an image [28], helping models capture spatial and semantic information without labelled data.
@@
In contrast, we present block masked self-supervised learning for hierarchical ViT pretraining, resulting in a model that is both powerful and efficient. To the best of our knowledge, no existing work jointly integrates hierarchical ViTs with block-level masked self-supervised learning for corneal confocal microscopy–based neuropathy diagnosis.
@@ 3.1 Overview of the Proposed Framework
    Hierarchical Multi-Scale Feature Extraction: HMSViT processes images through four hierarchical stages, starting with high-resolution inputs to capture fine details and gradually reducing spatial resolution to extract higher-level features. Instead of relying on complex modules (e.g., shifted windows or convolutions), our design uses a non-parametric pooling operator, with block-based local attention in early stages and global attention in deeper stages complemented by block-level absolute positional encodings.
@@ 3.2 Hierarchical Multi-Scale Vision Transformer (HMSViT)
The proposed method leverages the strengths of ViTs while addressing their limitations in dense medical imaging tasks specifically, for corneal nerve fibre segmentation and DPN diagnosis. By progressively extracting features at multiple scales, the network captures both fine-grained local details and broader global context, leading to enhanced segmentation performance and improved diagnostic accuracy (see Figure 2).
@@ 3.2.1 High-Resolution Hierarchical Representations
Figure 2: The architecture of the Hierarchical Multi-Scale Vision Transformer (HMSViT). An input image is processed through four hierarchical stages. Early stages use block-based local attention on high-resolution feature maps, while later stages switch to global attention after pooling reduces the spatial dimensions. This dual-attention approach balances computational efficiency with the ability to capture both local details and global context.
@@
where $L=\frac{H}{p}\times\frac{W}{p}$ is the number of patches, $p$ is the patch size, and $d$ is the feature dimension. $\frac{H}{p}\times\frac{W}{p}$ is the output feature size.
@@
A non-parametric max pooling operator is then applied to each block to downsample the representation. Max pooling is selected to preserve the fine-grained structural details typical of CCM, where nerve fibers appear as thin, high-intensity structures against a noisy background. Max pooling is theoretically ideal as it retains the strongest activation—representing the nerve signal—within a local block, whereas average pooling tends to smooth out these details, potentially obscuring thin nerve signals. Furthermore, compared to strided convolutions, max pooling is parameter-free, reducing model complexity. The pooling operation is defined as:
@@ 3.2.2 Dual Attention Mechanisms
Given that high-resolution features in the early stages result in a large number of tokens ($L$), applying global self-attention becomes computationally intensive due to its $O(L^{2})$ complexity. To mitigate this, HMSViT employs a dual attention strategy that adapts to the network’s changing hierarchical stages, balancing computational efficiency with feature extraction quality (see Figure 2). The standard scaled dot-product attention is defined as:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (4)
where $Q$, $K$, and $V$ are the query, key, and value matrices, and $d_{k}$ is the dimension of the keys. Our dual mechanism applies this operation at two different scopes:
  • Block-based Local Attention: In the early stages of the network (e.g., Stages 1 and 2), where feature maps are large, we apply attention locally within each block of tokens. For a given block $B^{l}$, the query, key, and value matrices ($Q,K,V$) are derived exclusively from the tokens within that block. This constrains the attention mechanism to focus on local, fine-grained features and significantly reduces computational cost, as the sequence length for each attention calculation is small (e.g., 16 tokens).
  • Global Attention: In the deeper stages (e.g., Stages 3 and 4), the feature maps have been downsampled, resulting in fewer tokens. At this point, the model switches to global attention. Here, the $Q,K,V$ matrices are derived from the entire sequence of tokens in the feature map. This allows the model to capture long-range dependencies and integrate high-level semantic information across the entire image, which is computationally feasible due to the reduced number of tokens.
This hybrid strategy allows HMSViT to efficiently process high-resolution inputs by first capturing local patterns with block-based attention and then modeling global context with global attention in later stages. This ensures a comprehensive feature representation without the prohibitive cost of applying global attention at all layers.
@@ 3.2.3 Absolute positional encodings
To preserve spatial relationships—crucial in medical imaging—we add positional encodings to the input data. While standard ViTs use absolute or relative embeddings (with hierarchical models often favoring relative ones for consistency across varying resolutions), relative embeddings require modifying the attention mechanism and add computational overhead. In our method, we add learnable absolute positional encodings at the block level:
@@
3.3 Block-masked self-supervised learning
Figure 3: An overview of the block-masked self-supervised learning framework. Patches are grouped into blocks, a portion of which are randomly masked. The HMSViT encoder processes the visible blocks and is trained to reconstruct the original input, forcing it to learn robust spatial representations.
The SSL approach builds on the recent advances in MAE [35], where a large portion of the input (typically 75–90%) is masked out, and the model is trained to reconstruct the missing parts. This strategy forces the network to learn rich spatial relationships, as it must infer the structure of the masked regions from the visible context. However, in a hierarchical Vision Transformer (ViT) with progressive downsampling, directly applying standard MAE can disrupt spatial coherence across scales. To address this, we propose a block-masking strategy that groups small patches into larger blocks of $16\times 16$ pixels (comprising $4\times 4$ patches) before masking. This is motivated by two key factors. First, masking tiny tokens tends to force the model to reconstruct trivial low-level textures, such as edges and noise. By contrast, masking larger blocks helps shift the learning signal from simply ”filling in pixel noise” to ”understanding the scene”, requiring the model to infer higher-level structure. Second, masking larger blocks improves computational efficiency, as it results in fewer mask decisions and fewer visible tokens for the decoder to process. As shown in Figure 3, we (1) group the patches into block units, (2) randomly mask blocks at the block level, and (3) ungroup them back into smaller patches for our HMSViT.